# ASSESS: A Semantic and Structural Evaluation Framework for Statement Similarity

### v1

@@ Abstract
Statement autoformalization, the automated translation of statements from natural language into formal languages, has seen significant advancements, yet the development of automated evaluation metrics remains limited. Existing metrics for formal statement similarity often fail to balance semantic and structural information. String-based approaches capture syntactic structure but ignore semantic meaning, whereas proof-based methods validate semantic equivalence but disregard structural nuances and, critically, provide no graded similarity score in the event of proof failure. To address these issues, we introduce ASSESS (A Semantic and Structural Evaluation Framework for Statement Similarity), which comprehensively integrates semantic and structural information to provide a continuous similarity score. Our framework first transforms formal statements into Operator Trees to capture their syntactic structure and then computes a similarity score using our novel TransTED (Transformation Tree Edit Distance) Similarity metric, which enhances traditional Tree Edit Distance by incorporating semantic awareness through transformations. For rigorous validation, we present EPLA (Evaluating Provability and Likeness for Autoformalization), a new benchmark of 524 expert-annotated formal statement pairs derived from miniF2F and ProofNet, with labels for both semantic provability and structural likeness. Experiments on EPLA demonstrate that TransTED Similarity outperforms existing methods, achieving state-of-the-art accuracy and the highest Kappa coefficient. The benchmark, and implementation code will be made public soon.
@@ 1 Introduction
Formal languages such as Isabelle (Paulson, 1994), HOL Light (Harrison, 1996), Coq (Barras et al., 1999), and Lean (De Moura et al., 2015; Moura & Ullrich, 2021) have recently gained prominence within the mathematical community for their capacity to rigorously verify proofs. Nevertheless, formalizing mathematical content is a labor-intensive process that demands substantial time, effort, and a profound familiarity with these specialized languages and their corresponding mathematical libraries, such as Mathlib (The mathlib Community, 2020). Consequently, the task of autoformalization (Szegedy, 2020), defined as translating theorem statements and proofs from natural language into their formal counterparts, has become an active area of research.
@@
While autoformalization has rapidly advanced, the methods for evaluating its output have not kept pace. Existing evaluation metrics can be unified under the framework of assigning a similarity score, and they are constrained by a fundamental trade-off between capturing semantic meaning and preserving structural information. String-based metrics such as BLEU (Papineni et al., 2002) rely on surface-level n-gram overlap. This makes them sensitive to inconsequential lexical variations but largely ignores underlying semantic content. Conversely, proof-based approaches (Li et al., 2024; Liu et al., 2025a) can guarantee semantic equivalence through formal proof. However, this method is brittle; it is fundamentally constrained by the capabilities of the underlying automated theorem prover, leading to a high false negative rate. Furthermore, when a proof cannot be found, these approaches fail to provide any meaningful similarity score. In addition, syntax-based metrics (Jiang et al., 2023) operate at the shallowest level of grammatical compliance, offering no semantic or structural insights. The LLM-as-a-Judge (Ying et al., 2024a) approach, while powerful, introduces prohibitive costs and significant reproducibility concerns. These collective shortcomings reveal a critical gap in the field: the need for an automated evaluation metric that robustly combines semantic and structural information for statement similarity in a reproducible and efficient manner.
@@
In this work, we propose ASSESS (A Semantic and Structural Evaluation Framework for Statement Similarity), a novel two-stage framework for evaluating formal statement pairs. The core of ASSESS is its novel metric, TransTED (Transformation Tree Edit Distance) Similarity. It is designed to be sensitive to both semantic and structural nuances while remaining computationally efficient, requiring only CPU resources. In the first stage, ASSESS leverages the Lean Language Server to parse each formal statement pair into its Operator Tree, a representation that captures deeper structural information than raw text. To overcome the semantic limitations of a standard Tree Edit Distance (TED) (Zhang & Shasha, 1989), the second stage augments the TED calculation with a curated set of transformations. This augmentation enables TransTED Similarity to robustly measure statement similarity where purely structural or semantic methods fail.
To enable a rigorous evaluation in the absence of a standard benchmark, we introduce EPLA (Evaluating Provability and Likeness for Autoformalization). We constructed this benchmark by first translating informal statements from the miniF2F-test (Zheng et al., 2022) and ProofNet-test (Azerbayev et al., 2023) datasets using the Herald Translator (Gao et al., 2024) and Gemini 2.5 Pro (Comanici et al., 2025). Successful translations were then paired with their ground-truth counterparts and subsequently annotated by human experts along two axes: semantic provability and structural likeness. The resulting benchmark comprises 524 annotated pairs, drawn from EPLA-miniF2F (373 pairs) and EPLA-ProofNet (151 pairs). To our knowledge, this represents the largest and most comprehensive dataset dedicated to this evaluation task.
@@
Experimental results demonstrate that our proposed TransTED Similarity metric consistently outperforms a range of baselines, establishing a new state-of-the-art in both accuracy (78.82% and 70.86%) and the Cohen’s Kappa coefficient (0.46 and 0.40) across both the EPLA-miniF2F and EPLA-ProofNet datasets. Our metric provides a more balanced assessment than brittle proof-based methods, is more efficient and reproducible than LLM-based judges, and exhibits superior robustness compared to n-gram-based metrics. Finally, a detailed ablation study confirms that our novel transformation component is the critical factor behind these performance gains, validating its effectiveness in capturing semantic equivalence. This highlights its success in providing a more balanced and reliable measure of statement similarity.
@@
    We propose a novel method that leverages the Lean Language Server to automatically parse formal statements into Operator Trees. This representation enables the direct application of Tree Edit Distance (TED) similarity for evaluating statement similarity.
@@
    We introduce TransTED Similarity, a semantically-aware metric that extends TED similarity by incorporating transformations directly into the distance calculation. TransTED similarity provides a more holistic similarity score, establishing a new state-of-the-art.
@@
    We present EPLA, the first large-scale and most comprehensive benchmark for this evaluation task, comprising 524 formal statement pairs meticulously annotated by domain experts, to facilitate rigorous and reproducible evaluation.
@@ 2 Related Work
The goal of autoformalization is to translate informal mathematics into formally verified code. Research in this area, particularly for theorem statements, has evolved significantly. While early work often relied on neural machine translation (Wang et al., 2018; Cunningham et al., 2023), the transformative impact of Large Language Models (LLMs) has given rise to three dominant strategies. These include exploring the efficacy of few-shot prompting (Wu et al., 2022; Agrawal et al., 2022; Zhou et al., 2024); improving capabilities by fine-tuning LLMs (Gao et al., 2024; Lu et al., 2024b; Liu et al., 2025b; Wu et al., 2025) on relevant parallel statements; and integrating retrieval-augmented generation (Zhang et al., 2024) to achieve enhanced performance.
@@
Automated evaluation in the context of autoformalization denotes methods and metrics that assign quantitative scores to machine-generated formal statements, with the goal of estimating their similarity relative to a natural-language source (Weng et al., 2025). In the realm of automated evaluation, early efforts predominantly employ metrics based on grammatical validity (Jiang et al., 2023) and string similarity (Azerbayev et al., 2023), yet these struggle with semantic understanding. FormalAlign (Lu et al., 2024a) innovatively integrates autoformalization with evaluation by simultaneously generating a formal statement and its corresponding evaluation score. While this represents a significant contribution to the fields of autoformalization and automated evaluation, its scoring mechanism cannot be used as a standalone evaluation metric. Simultaneously, cross-provability (Murphy et al., 2024; Li et al., 2024; Liu et al., 2025a; Poiroux et al., 2024) between formal statements emerges as a widely accepted standard for automated evaluation, but its effectiveness is constrained by the current progress in automated theorem proving.
@@
Operator Trees (OPTs) represent mathematical expressions as syntax trees, with operators as internal nodes and operands as leaves (Zanibbi et al., 2002). Compared with sequence-based formula representations, OPTs explicitly capture the hierarchical structure and semantic relations within expressions, preserving operator precedence and operand dependencies (Zanibbi & Blostein, 2012; Hu et al., 2013). These properties have made OPTs foundational to applications in mathematical information retrieval (MIR). For instance, systems extract structural features from OPTs for similarity searching (Zhong & Zanibbi, 2019) or combine them with other representations for formula retrieval, as in the MCAT system (Kristianto et al., 2016). The utility of OPTs also extends to deep learning, where models like MathBERT (Peng et al., 2021) integrate OPT structures during pretraining to enhance semantic understanding , and encoders such as FORTE (Wang et al., 2021) learn formula representations directly from OPT-based inputs.
Figure 1: Illustration of ASSESS. (a) EPLA Benchmark: The dataset is created by using translators to generate formal pairs, which are processed through compilation and human evaluation. (b) TED Similarity: The baseline TED Similarity is computed by converting an FL pair into Operator Trees, calculating and normalizing the TED. (c) TransTED Similarity: The novel TransTED Similarity first merges an FL pair into an equality. It then employs a transformation loop, guided by a search that uses TED as heuristic function, to find and normalize the minimal achievable distance.
@@ 3 Methodology
This section details our proposed framework ASSESS, which is illustrated in Figure 1. We first establish the theoretical foundations in Section 3.1. We then present the two core metrics: an initial evaluation using a standard Tree Edit Distance (TED) Similarity (Section 3.2), followed by our primary contribution, the semantically-enhanced TransTED Similarity metric (Section 3.3).
@@ 3.1 Preliminaries
Crucially, a pseudometric differs from a metric in that $d(x,y)=0$ does not necessarily imply $x=y$. This property is essential for our application, as structurally distinct formal statements can be semantically equivalent (e.g., $a+b$ and $b+a$). For generality, we allow the codomain to be the extended non-negative real numbers $0,\infty]$, noting that the operation $\infty$-$\infty$ is undefined.
@@
Let $G=(V,E,w)$ be an undirected, weighted graph, where $V$ is the set of vertices, $E$ is the set of edges, and $w:E\to[0,\infty]$ is the edge weight function. The shortest path distance $d:V\times V\to[0,\infty]$ is defined as:
@@
By convention, $d(v,v)=0$. If no path with finitely many edges exists between $u$ and $v$, the set of paths is empty, and the infimum over an empty set is $\infty$, so $d(u,v)=\infty$. This function $d$ satisfies the required axioms, confirming that $(V,d)$ constitutes a pseudometric space.
@@ 3.2.1 OPT Construction
To capture the hierarchical structure of formal statements, we represent each as a labeled, ordered Operator Tree (OPT). We leverage the Lean Language Server to parse a statement, from which we derive the tree’s topology based on the nested scopes of its elements: operators become parent nodes and their arguments become ordered children. During construction, we apply two additional transformations to standardize the tree structure:
@@ 3.2.2 Metric Calculation
With formal statements converted to OPTs, we can instantiate the pseudometric space defined in Section 3.1. The set of vertices $X$ becomes the set of all possible OPTs, and the distance function $d$ is the Tree Edit Distance (TED).
@@
  • Delete a node, connecting its children to its parent in order. Cost: $c_{\text{del}}$.
  • Insert a node, the inverse of a delete operation. Cost: $c_{\text{ins}}$.
  • Relabel a node by changing its label. Cost: $c_{\text{rel}}$.
@@
To enable meaningful comparisons across trees of varying sizes, we define TED Similarity by normalizing the absolute Tree Edit Distance (TED) relative to tree size.
@@
This formulation assumes unit costs for all operations ($c_{\text{del}}=c_{\text{ins}}=c_{\text{rel}}=1$), a standard convention we adopt (Zhang & Shasha, 1989). The denominator $\max\{|T_{1}|,|T_{2}|\}$ serves as a normalization factor, representing the cost of deleting all nodes in the larger tree. This ensures the similarity score is bounded in $[0,1]$, where 1 indicates identical trees.
@@ 3.3 TransTED Similarity
While TED Similarity effectively captures structural differences, it is syntactically rigid and penalizes semantically equivalent expressions, such as $a+b$ and $b+a$. To address this, we introduce TransTED Similarity, which augments the TED with semantic awareness through transformations.
@@ 3.3.1 Theoretical Framework
In our framework, we define a concept, transformation, to integrate semantic awareness into distance functions. We start with an example: $i=j$ implies that $f(i)=f(j)$. In other words, $f(i)=f(j)$ is no stronger than $i=j$, because to show $f(i)=f(j)$ suffices to show $i=j$. We say that the comparison of $f(i)$ and $f(j)$ can be transformed to the comparison of $i$ and $j$. Generally, a weaker comparison can be transformed to a stronger one. Therefore, we hope that the distance between OPTs of the transformed comparison is no smaller than that between OPTs of the original one.
Formally, let $X$ be the set of all Operator Trees (OPTs). We define a new pseudometric, $d^{*}$, on $X$ that is constrained by the following two properties:
@@
  • Monotonicity under Transformation: If a pair of formal expressions $(e_{x},e_{y})$ can be transformed into a logically stronger pair $(e_{u},e_{v})$ (i.e., if $e_{u}=e_{v}$ implies $e_{x}=e_{y}$), then the metric must satisfy $d^{*}(\operatorname{OPT}(e_{x}),\operatorname{OPT}(e_{y}))\leq d^{*}(\operatorname{OPT}(e_{u}),\operatorname{OPT}(e_{v}))$.
Finding the largest pseudometric $d^{*}$ that satisfies these conditions can be formulated as a linear programming problem. The following theorem establishes that this problem has a unique optimal solution, which we define as TransTED. The full proof is provided in Appendix  A.
@@
Let $X$ be an arbitrary set, $b:X\times X\to[0,\infty]$ be a function and $T$ is a subset of $(X\times X)^{2}$. Consider the function space
$\mathcal{F}:=\left\{f\Bigg|f\text{ is a pseudometric on }X,\text{ and }\begin{cases}\forall x,y\in X,f(x,y)\leq b(x,y),\\ \forall((x,y),(u,v))\in T,f(x,y)\leq f(u,v).\end{cases}\right\}.$

### v2

@@ Abstract
Despite significant strides in statement autoformalization, a critical gap remains in the development of automated evaluation metrics capable of assessing formal translation quality. Existing metrics often fail to balance semantic and structural information: string-based methods neglect semantics, whereas proof-based approaches offer no graded similarity when proofs fail. To address these issues, we introduce ASSESS (A Semantic and Structural Evaluation Framework for Statement Similarity), which captures syntactic structure by transforming formal statements into operator trees and computes a real-valued similarity score using our novel TransTED (Transformation Tree Edit Distance) Similarity metric by incorporating semantic transformations. For rigorous validation, we present EPLA (Evaluating Provability and Likeness for Autoformalization), a benchmark comprising 1,247 expert-annotated formal statement pairs derived from miniF2F and ProofNet, distinctively labeled for both semantic provability and structural likeness. Experiments on the EPLA benchmark demonstrate that TransTED Similarity surpasses existing methods, achieving state-of-the-art accuracy and Kappa score. The benchmark dataset, code, and detailed experimental results are available at https://github.com/XiaoyangLiu-sjtu/ASSESS.
@@ 1 Introduction
Formal languages such as Isabelle (Paulson, 1994), HOL Light (Harrison, 1996), Rocq (Barras et al., 1997), and Lean (de Moura et al., 2015; Moura and Ullrich, 2021) have recently gained prominence within the mathematical community for their capacity to rigorously verify proofs. Nevertheless, formalizing mathematical content is a labor-intensive process that demands substantial time, effort, and a profound familiarity with these specialized languages and their corresponding mathematical libraries, such as Mathlib (mathlib Community, 2020). Consequently, the task of autoformalization (Szegedy, 2020), defined as translating theorem statements and proofs from natural language into their formal counterparts, has become an active area of research.
@@
While autoformalization has advanced rapidly, methods for evaluating its output have lagged behind. Existing metrics, which generally function by assigning a similarity score, are constrained by a fundamental trade-off between capturing semantic meaning and preserving structural information. String-based metrics such as BLEU (Papineni et al., 2002) rely on surface-level n-gram overlap, rendering them sensitive to inconsequential lexical variations while remaining blind to underlying semantic content. Conversely, proof-based approaches (Li et al., 2024; Liu et al., 2025a) can guarantee semantic equivalence but are limited by prover brittleness, resulting in high false negatives and no graded feedback for unproven statements. Additionally, syntax-based metrics (Jiang et al., 2023) operate merely at the level of grammatical compliance, offering negligible semantic or structural insights. Finally, the LLM-as-a-Judge (Ying et al., 2024a) approach, while powerful, introduces prohibitive costs and significant reproducibility concerns. These collective shortcomings highlight a critical gap: the need for an automated evaluation metric that robustly integrates semantic and structural information to assess statement similarity in a reproducible and efficient manner.
@@
In this work, we propose ASSESS (A Semantic and Structural Evaluation Framework for Statement Similarity), a novel two-stage framework for evaluating formal statement pairs. The core of ASSESS is its novel metric, TransTED (Transformation Tree Edit Distance) Similarity. It is designed to capture both semantic and structural nuances, offering a reproducible solution that relies exclusively on CPU resources. In the first stage, ASSESS leverages the Lean Language Server to parse each formal statement pair into their operator trees (OPTs), a representation that captures hierarchical structure more effectively than raw text. To overcome the semantic limitations of a standard Tree Edit Distance (TED) (Zhang and Shasha, 1989), the second stage augments the TED with a curated set of transformations. This augmentation enables TransTED Similarity to robustly measure statement similarity where purely structural or semantic methods fail.
To enable rigorous evaluation, we introduce EPLA (Evaluating Provability and Likeness for Autoformalization). We constructed this benchmark by autoformalizing informal statements from the miniF2F-test (Zheng et al., 2022) and ProofNet-test (Azerbayev et al., 2023) datasets, utilizing a combination of two domain-specific and two general-purpose models. We filtered the candidate formalizations using the Lean compiler to ensure syntactic validity. The surviving instances were paired with ground-truth references and expert-annotated based on semantic provability and structural likeness. The benchmark contains 1,247 annotated pairs in total, with 831 from EPLA-miniF2F and 416 from EPLA-ProofNet.
@@
Experimental results demonstrate that TransTED Similarity consistently outperforms all baselines. It establishes a new state-of-the-art by achieving 70.16% accuracy and a 0.35 Kappa score on EPLA-miniF2F, alongside 67.31% accuracy and a 0.30 Kappa score on EPLA-ProofNet. Compared to existing approaches, our metric provides a more robust assessment than brittle proof-based methods, is more reproducible than LLM-based judges, and surpasses the performance of n-gram-based metrics. Furthermore, a detailed ablation study identifies the transformation component as the key factor in these performance gains. This confirms that our mechanism bridges the gap between semantic equivalence and structural likeness by dynamically aligning syntactically diverse yet logically identical expressions.
@@
    We propose a novel method that leverages the Lean Language Server to automatically parse formal statements into operator trees. This structured representation enables the direct application of TED Similarity to quantify statement similarity.
@@
    We introduce TransTED Similarity, which augments the traditional TED Similarity by incorporating semantic transformations into the distance computation. This semantic-aware metric achieves state-of-the-art performance on the EPLA benchmark.
@@
    We present EPLA, a comprehensive benchmark tailored for this evaluation task, comprising 1,247 formal statement pairs annotated by domain experts.
@@ 2 Related Work
The goal of autoformalization is to translate informal mathematics into formal code. Research in this area, particularly for theorem statements, has evolved significantly. While early work often relied on neural machine translation (Wang et al., 2018; Cunningham et al., 2022), the transformative impact of Large Language Models (LLMs) has given rise to three dominant strategies. These include exploring the efficacy of few-shot prompting (Wu et al., 2022; Agrawal et al., 2022; Zhou et al., 2024); improving capabilities by fine-tuning LLMs (Gao et al., 2025; Lu et al., 2024; Liu et al., 2025b; Wu et al., 2025; Yu et al., 2025) on relevant parallel statements; and integrating retrieval-augmented generation (Zhang et al., 2024) to achieve enhanced performance.
@@
Automated evaluation in the context of autoformalization denotes metrics that assign quantitative scores to machine-generated formal statements, with the goal of estimating their similarity relative to a natural-language source (Weng et al., 2025). Early efforts relied on grammatical validity (Jiang et al., 2023) and string similarity (Azerbayev et al., 2023), yet these often struggle with semantic understanding. While FormalAlign (Lu et al., 2025) integrates evaluation with autoformalization, its internal scoring mechanism cannot serve as a standalone metric. Simultaneously, cross-provability (Murphy et al., 2024; Li et al., 2024; Liu et al., 2025a; Poiroux et al., 2025) between formal statements emerges as a widely accepted standard for automated evaluation, but its effectiveness is constrained by the current progress in automated theorem proving. GTED (Liu et al., 2025c) pioneered the use of operator trees for this evaluation task; however, it suffered from implementation instabilities and its transformation mechanism was confined to simple variable renaming. In contrast, our framework formalizes these concepts into a rigorous pseudometric space and introduces proof-based transformations to capture deep semantic alignment.
@@
Operator trees (OPTs) represent mathematical expressions as syntax trees, with operators as internal nodes and operands as leaves (Zanibbi et al., 2002). Compared with sequence-based formula representations, OPTs explicitly capture the hierarchical structure and semantic relations within expressions, preserving operator precedence and operand dependencies (Zanibbi and Blostein, 2012; Hu et al., 2013). These properties have made OPTs foundational to applications in mathematical information retrieval (MIR). For instance, systems extract structural features from OPTs for similarity searching (Zhong and Zanibbi, 2019) or combine them with other representations for formula retrieval, as in the MCAT system (Kristianto et al., 2016). The utility of OPTs also extends to deep learning, where models like MathBERT (Peng et al., 2021) integrate OPT structures during pretraining to enhance semantic understanding , and encoders such as FORTE (Wang et al., 2021) learn formula representations directly from OPT-based inputs.
Figure 1: Overview of the ASSESS Framework. (a) EPLA Benchmark: This dataset is constructed using four distinct translators to generate Formal Language (FL) pairs, followed by compilation checks and human evaluation. (b) TED Similarity: A baseline metric computed by converting FL pairs into operator trees to calculate the Tree Edit Distance (TED) similarity. (c) TransTED Similarity: A novel metric that reformulates an FL pair as an equality. It conducts a tree search guided by specific tactic commands, utilizing TED as a heuristic to minimize the edit distance. The search halts upon satisfying any of the following conditions: successful proof generation (Proved), Node Limit Exceeded (NLE), or Time Limit Exceeded (TLE).
@@ 3 Methodology
This section details our proposed framework, ASSESS, as illustrated in Figure 1. We begin by establishing the theoretical foundations in Section 3.1. Subsequently, we present the two core metrics: the basic TED Similarity in Section 3.2, followed by our primary contribution, the semantically-aware TransTED Similarity in Section 3.3.
@@ 3.1 Preliminaries
Crucially, a pseudometric differs from a metric in that $d(x,y)=0$ does not necessarily imply $x=y$. This property is essential for our application, as structurally distinct formal statements can be semantically equivalent (e.g., $a+b$ and $b+a$).
@@
Let $G=(V,E,w)$ be an undirected, weighted graph, where $V$ is the set of vertices, $E$ is the set of edges, and $w:E\to\mathbb{R}_{\geq 0}$ is the edge weight function. The shortest path distance $d:V\times V\to\mathbb{R}_{\geq 0}$ is defined as:
@@
By convention, $d(v,v)=0$. This function $d$ satisfies the required axioms, confirming that $(V,d)$ constitutes a pseudometric space.
@@ 3.2.1 OPT Construction
To capture the hierarchical structure of formal statements, we represent each as a labeled, ordered operator tree. We leverage the Lean Language Server to parse a formal statement, from which we derive the tree’s topology based on the nested scopes of its elements: operators become parent nodes and their arguments become ordered children. During construction, we apply two additional transformations to standardize the tree structure:
@@ 3.2.2 Metric Calculation
With formal statements converted to OPTs, we can instantiate the pseudometric space defined in Section 3.1. The set of vertices $X$ becomes the set of all possible OPTs, and the distance function $d$ is the Tree Edit Distance.
@@
  • Deleting a node, connecting its children to its parent in order. Cost: $c_{\text{del}}$.
  • Inserting a node, the inverse of a delete operation. Cost: $c_{\text{ins}}$.
  • Relabelling a node by changing its label. Cost: $c_{\text{rel}}$.
@@
To enable meaningful comparisons across trees of varying sizes, we define TED Similarity by normalizing the absolute Tree Edit Distance relative to tree size.
@@
This formulation assumes unit costs for all operations ($c_{\text{del}}=c_{\text{ins}}=c_{\text{rel}}=1$), a standard convention we adopt (Zhang and Shasha, 1989). The denominator $\max\{|T_{1}|,|T_{2}|\}$ serves as a normalization factor, representing the cost of deleting all nodes in the larger tree.
@@ 3.3 TransTED Similarity
While TED Similarity effectively captures structural differences, it is syntactically rigid and penalizes semantically equivalent expressions, such as $a+b$ and $b+a$. We therefore introduce TransTED Similarity, which integrates semantic awareness into the distance computation via transformations.
@@ 3.3.1 Theoretical Framework
In our framework, we define a concept, transformation, to integrate semantic awareness into distance functions. We start with an example: $i=j$ implies that $f(i)=f(j)$. In other words, $f(i)=f(j)$ is no stronger than $i=j$. We say that the contrast of $f(i)$ and $f(j)$ can be transformed to the contrast of $i$ and $j$. Generally, a weaker contrast can be transformed to a stronger one. We notice that the two sides of a weak contrast are usually near in semantic. Therefore the distance between the original contrast should be no larger than that between the transformed one. For compatibility with TED, we will define a new distance between OPTs instead of expressions.
Here we provide a formal definition of TransTED. Let $X$ be the set of all operator trees. We define a new pseudometric $d^{*}$ on $X$ that is constrained by the following two properties:
@@
  • Monotonicity under transformation: If a pair of formal expressions $(e_{x},e_{y})$ can be transformed into a logically stronger pair $(e_{u},e_{v})$ (i.e., if $e_{u}=e_{v}$ implies $e_{x}=e_{y}$), then the metric must satisfy $d^{*}(\text{{OPT}}(e_{x}),\text{{OPT}}(e_{y}))\leq d^{*}(\text{{OPT}}(e_{u}),\text{{OPT}}(e_{v}))$.
Finding the largest pseudometric $d^{*}$ that satisfies these conditions can be formulated as a linear programming problem. The following theorem establishes that this problem has a unique optimal solution, which we define as TransTED. The full proof is provided in Appendix  C.
@@
Let $X$ be an arbitrary set, $b:X\times X\to\mathbb{R}_{\geq 0}$ be a function and $T$ is a subset of $(X\times X)^{2}$. Consider the function space
$\mathcal{F}:=\left\{f\Bigg|f\text{ is a pseudometric on }X,\text{ and }\begin{cases}\forall x,y\in X,f(x,y)\leq b(x,y),\\ \forall((x,y),(u,v))\in T,f(x,y)\leq f(u,v)\end{cases}\right\}.$